A global financial institution transfers a large data file every evening from offices around the world to its London headquarters. Once the file is recieved, it must be cleaned and partitioned before being stored in a company's data warehouse. Each file is the same size and the time required to transfer, clean, and partition a file is normally distributed, with a mean of 1.5 hours and standard deviation of 15 minutes. If a manager must be present until 85% of the files are transferred, cleaned, and partitioned, how long will the manager need to be there?


1
Expert's answer
2021-06-24T12:31:39-0400

Let "X=" the time required to transfer: "X\sim N(\mu, \sigma^2)."

Given "\mu=90, \sigma=15."



"P(X<x)=P(Z<\dfrac{x-90}{15})=0.85"

"\dfrac{x-90}{15}\approx1.036433"

"x=105.5\ min"